# Unsupervised Learning of Effective Quantum Impurity Models

### v2

@@ V Comparison of models
Figure 7: Comparison of performance metrics for the various optimized effective models considered in this work, with the AIM as the benchmark bare model. Results obtained by NRG. (a) Kondo temperature $T_{K}$ as a function of $U_{d}$, comparing the AIM (points) and $\mathcal{F}$-learned Kondo result (blue line) with the UML-optimized effective models: minimal (green), MC (orange), and extended MC (red). Corresponding $T_{K}$ relative error shown in (b) and normalized observables shown in (c). (d) Impurity contribution to the entropy $S_{\rm imp}(T)$ vs $T/U_{d}$ for the bare AIM with $U_{d}=0.4$ (points) and the optimized effective models (lines). In all cases we set $0.38V^{2}/U_{d}=$ 0.3.
In this section we consider the relative performance of the various effective models introduced in this paper, optimized by UML to best describe the the physics of the AIM as our benchmark bare model. As we have shown, it is in general not possible to simultaneously capture within a simplified effective model the universal low-temperature thermodynamics (as characterized by the Kondo temperature $T_{K}$) as well as the behavior of static thermal expectation values of local operators. Any effective model must therefore strike a compromise in reproducing these features. Different methods may target different aspects of the physics and be better suited to different applications.
@@ VI Conclusion
In this paper, we focused on developing new ML-inspired methodologies for the automatic derivation of expressive and accurate effective models for quantum impurity systems. To establish the feasibility and performance of our UML method, we studied in detail the AIM as the bare model. This is because the AIM is a very well-known paradigm [36], which we can solve exactly down to $T=0$ using NRG [52] for benchmarking purposes. The same approach can in principle be used for more complicated bare model systems as described in Sec. III.2.4. However, we recognize that useful applications of UML may require integration with ab initio methods to treat realistic systems [74, 75, 76, 77, 73, 4]. This is beyond the scope of the current work, but we discuss possible ways to do this as an outlook in Appendix D.
@@
Care must be taken however, since thermodynamic observables are not invariant under RG and so bare an effective models with different RG flows may have different static expectation values. Thus, the naive matching of observables does not imply that low-energy scales, such as the Kondo temperature, are correctly recovered. We confirm this directly, already at the simplest possible example of the Anderson-Kondo model mapping. To mitigate this we identify and construct the best observables to match, based on an analysis of their RG relevance. We show that this can be done while maintaining convexity of the optimization, and that low-energy scales can be recovered. The accuracy of our effective models can be systematically improved by increasing the complexity of the effective impurity. Although this clearly offers no advantage at the level of the AIM, the approach is general and systematic, meaning that complex systems beyond reach of existing methods can be boiled down to tractable effective models.

### v3

@@ V Comparison of models
In this section we consider the relative performance of the various effective models introduced in this paper. Each is optimized by UML to capture the physics of the AIM, which serves as our benchmark bare model. As we have shown, it is in general not possible to simultaneously capture within a simplified effective model the universal low-temperature thermodynamics (as characterized by the Kondo temperature $T_{K}$) as well as the behavior of static thermal expectation values of local operators. Any effective model must therefore strike a compromise in reproducing these features. Different methods may target different aspects of the physics and be better suited to different applications.
@@
Figure 8: Comparison of performance metrics for the optimized effective models considered in this work, taking the AIM as the bare model. (a) Kondo temperature $T_{K}$ vs $U_{d}$, comparing the AIM (points) and $\mathcal{F}$-learned Kondo result (blue line) with the UML-optimized effective models: minimal (green), MC (orange), and extended MC (red). Corresponding $T_{K}$ relative error shown in (b). Observables $\langle\hat{O}\rangle$ normalized by their $U_{d}\to\infty$ values shown in panel (c), with $\hat{O}=\mathbf{S}_{d}\cdot\mathbf{S}_{0}$ (solid lines, top) and $(\hat{n}_{0}-1)^{2}$ (dashed lines, bottom). (d) Entropy $S_{\rm imp}(T)$ vs $T/U_{d}$ for the bare AIM ($U_{d}=0.4$, points) and the optimized effective models (lines). We set $0.38V^{2}/U_{d}=$ 0.3 throughout.
@@ VI Conclusion
In this paper, we focused on developing new ML-inspired methodologies for the automatic derivation of expressive and accurate effective models for quantum impurity systems. To establish the feasibility and performance of our UML method, we studied in detail the AIM as the bare model. This is because the AIM is a very well-known paradigm [36], which we can solve exactly down to $T=0$ using NRG [55] for benchmarking purposes. The same approach can in principle be used for more complicated bare model systems as described in Sec. III.2.5. However, we recognize that useful applications of UML may require integration with ab initio methods to treat realistic systems [77, 78, 79, 80, 76, 4]. This is beyond the scope of the current work, but we discuss possible ways to do this as an outlook in Appendix D.
@@
Care must be taken however, since thermodynamic observables are not invariant under RG and so bare an effective models with different RG flows may have different static expectation values. Thus, the naive matching of observables does not imply that low-energy scales, such as the Kondo temperature, are correctly recovered. We confirm this directly, already at the simplest possible example of the Anderson-Kondo model mapping. To mitigate this we identify and construct the best observables to match, based on an analysis of their RG relevance. We show that this can be done while maintaining convexity of the optimization, and that low-energy scales can be recovered. The accuracy of our effective models can be systematically improved by increasing the complexity of the effective impurity. Although this clearly offers no advantage at the level of the AIM, UML produces dramatically simplified effective models when applied to more complex systems, consisting of 3 or more interacting active orbitals. The accuracy and efficiency of UML may be further improved by formulating bare and effective models in the ‘Natural Orbital’ (NO) basis [86]. Since entanglement between NOs and the rest of the fermionic bath is small by construction, local observables to be matched in UML should flow weakly, which we have showed is a desirable property when formulating effective models.